OpenStudy (anonymous):

the length of a rectangular frame is 6 cm more than 9 times their sum. find the integers.

OpenStudy (anonymous):

???? Pls check if u hv typed the question as it is given in book or have u missed something....

OpenStudy (anonymous):

oh wow, yes i did. the length of a rectangular frame is 6cm more than the width. the area inside the frome is 55sq cm. find width of frame.

OpenStudy (anonymous):

Let width = x cm then length = (x + 6) cm area of rectangle = length * width ATQ 55 = (x + 6)(x) 55 = x^2 + 6x x^2 + 6x - 55 = 0 x^2 + 11x - 5x - 55 = 0 x(x + 11) -5(x + 11) = 0 (x - 5)(x + 11) = 0 if x - 5 = 0, then x = 5 if x + 11 = 0, then x = -11 but length can never be negative, so x = 5 is the right value so width = x cm = 5 cm length = (x+6) cm = (5+6) cm = 11 cm